OpenStudy (anonymous):

A car traveling north at 40 mph and a truck traveling east at 30 mph leave an intersection at the same time. At what rate will the distance between them be changing 3 hours later? (Let D(t) denote the distance between them at time t .)

OpenStudy (anonymous):



OpenStudy (anonymous):

so do i differentiate pythagorean theorem?

OpenStudy (anonymous):

\[D^2=x^2+y^2\] \[2DD'=2xx'+2yy'\] \[x'=40, y'=30\]

OpenStudy (anonymous):

ok makes sense so far

OpenStudy (anonymous):

yeah differentiate using pythagoras now one one hour later x and y are 30 and 40 respectively, so you know by pythagoras that D =50 plug in the numbers, solve for D'

OpenStudy (anonymous):

oh that was a mistake, 3 hours later

OpenStudy (anonymous):

so x and y are 90 and 120

OpenStudy (anonymous):

3 hours later x is 120, and y is 90 so by pythagoras D is 150

OpenStudy (anonymous):

how did you get 150

OpenStudy (anonymous):

would it not be 210

OpenStudy (anonymous):

hmm maybe

OpenStudy (anonymous):

or sort of 210

OpenStudy (anonymous):

you have a 3 - 4 - 5 right triangle

OpenStudy (anonymous):

i meant sqrt 210

OpenStudy (anonymous):

3- 4- 5 30 -40 -50 90- 120 - 150

OpenStudy (anonymous):

it is 150

OpenStudy (anonymous):

whoa that just blew my mind what is that all about?

OpenStudy (anonymous):

ok i see you just multiplied all by 30 right

OpenStudy (anonymous):

\[\sqrt{90^2+120^2}=\sqrt{22500}\]

OpenStudy (anonymous):

right exactly just used the ratios, but you get 150 in any case i am sure

OpenStudy (anonymous):

so 150 is the answer thats my rate at which the distance is changing

OpenStudy (anonymous):

oh no , that is D

OpenStudy (anonymous):

you have \[2CC'=2xx'+2yy'\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[2D'D=2xx'+2yy'\]

OpenStudy (anonymous):

or better yet \[DD'=xx'+yy'\]

OpenStudy (anonymous):

you know \[x'=40,x=120, y'=30,y=90, D=150\] and you want D'

OpenStudy (anonymous):

plug in the numbers, solve for D'

OpenStudy (anonymous):

\[150D'=120\times 40+90\times 30\] etc

OpenStudy (anonymous):

ok, i guess i understand it now for this problem but the real problem I'm having is finding what my variables to assign in each problem

OpenStudy (anonymous):

you get to pick the variable. trick is to find the equation relating them

OpenStudy (anonymous):

and i just worked that out and got 50 and that is not correct

OpenStudy (anonymous):

look for similar triangles, pythagoras etc ok lets see

OpenStudy (anonymous):

\[150D'=40\times 120+30\times 90\] yeah i get 50 as well

OpenStudy (anonymous):

why is 50 mph wrong? it looks good to me

OpenStudy (anonymous):

beats me man... this stuff is mind boggling to me. I've actually posted like 3 or 4 questions up here from my homework this week and no one has yet been able to help me...

OpenStudy (anonymous):

how do you know it is wrong?

OpenStudy (anonymous):

i am fairly sure it is right

OpenStudy (anonymous):

ok nevermind man it was asking for it in kilometers i don't know why... so i converted it and it was correct

OpenStudy (anonymous):

lord just to confuse you ok fine, now we got it right?